OpenStudy (anonymous):

Solve 3x2 + 4x = 2. TEACH me how to solve this.. I know what to do once I have it in the right forum... but I do not understand how to simplify it down. a. quantity of 2 plus or minus square root of 10 all over 6 b. quantity of negative 2 plus or minus 2 square root of 10 all over 3 c. quantity of negative 2 plus or minus square root of 10 all over 3 d. quantity of 4 plus or minus square root of 10 all over 3

OpenStudy (radar):

\[3x^{2}+4x-2=0\]A=3, B=4 and C=-2

OpenStudy (radar):

The quadratic formula for solving for x is:\[-B \pm \sqrt{B ^{2}-4AC}\over2A\] Is that enough to get you started?

OpenStudy (radar):

Pllug in the values for A, B, and C and see what happens.

OpenStudy (anonymous):

the plugging in part i do get, but how does A become 3.. do we just totaly ignore the x?

OpenStudy (radar):

\[-4\pm \sqrt{16-(4)(3)(-2)}\over6\]\[-4\pm \sqrt{40}\over6\]\[-4\pm2\sqrt{10}\over6\]\[-2\pm \sqrt{10}\over3\]

OpenStudy (radar):

Now which of the sentences say that?

OpenStudy (radar):

Did you follow the plug in, and subsequent algebra to the final solution?

OpenStudy (anonymous):

3x2+4x−2=0 A=3, B=4 and C=-2 ^does the x get ignored?

OpenStudy (radar):

No x is what the results will be, x= -B+/sqrt(-B^2-4AC)/2A

OpenStudy (radar):

That is the solution for x.

OpenStudy (anonymous):

so 3=a i dont have to simplify it?

OpenStudy (radar):

\[x=(-B \pm \sqrt{B ^{2}-4AC})/2A\]

OpenStudy (radar):

My fifth post is the soluition for x using the quadratic. Just imagine to the left of each line there is a "x="

OpenStudy (radar):

A, B, are the coefficients of the variable and C is the constant.

OpenStudy (radar):

A=3 (3x^2) B=4 (4x) C=-2 -2

OpenStudy (radar):

Simplify where needed for example\[\sqrt{40}=\sqrt{4}\sqrt{10}=2\sqrt{10}\]

OpenStudy (radar):

Do you follow joemost?

OpenStudy (radar):

Read sentence c. and look at my solution.

OpenStudy (radar):